Attempts: Score: /10 Attention: Due to a bug in Google Chrome, this page may not function correctly. lick here to learn more 4. Gravetter/Wallnau/Forzano, Essentials- Chapter 7 -End-of-chapter question 13 A population has a mean of μ-30 and a standard deviation of σ-8. Standard Normal Distribution Mean-0.0 Standard Deviation1.0 6179 3821 3 -2 0 0.30 32 for a If the population distribution is normal, what is the probability of obtaining a sample mean greater than M sample of n 4? 0 0.3085 ๏ 0.40 13 O cannot be determined 0 0.0228 O Type here to search

If the population distribution is positively skewed, what is the probability of obtaining a sample mean greater than M -32 for a sample of n 47 O 0.4013 O Cannot be determined O 0.0228 O 0.3085 If the population distribution is normal, what is the p sample of n 64? robability of obtaining a sample mean greater than M 32 for a O 0.0228 O 0.3085 O Cannot be determined O 0.4013 If the population distribution is positively skewed, what is the probability of obtaining a sample mean greater than M 32 for a sample of n 64? 0.3085 O 0.0228 0.4013 O cannot be determined
Answer #1

4 i)P(M>32)=P(Z>(32-30)*sqrt(4)/8)=P(Z>0.5)=0.3085

ii) can not be determined as we can not apply normal approximation on above

iii) P(M>32)=P(Z>(32-30)*sqrt(64)/8)=P(Z>2)=0.0228

iv) P(M>32)=P(Z>(32-30)*sqrt(64)/8)=P(Z>2)=0.0228